OpenStudy (anonymous):

Solve the logarithmic equation. Please show all of your work. log6(x+2)+log6(x-3)=1

OpenStudy (mertsj):

\[\log_{6}(x+2)(x-3)=1 \]

OpenStudy (mertsj):

\[6^{1}=(x+2)(x-3)\]

OpenStudy (mertsj):

\[x^2-x-6=6\]

OpenStudy (anonymous):

log6(x+2)+log6(x-3) = 1 by laws of logs log6 (x+2(x-3) = 1 (x+2)(x-3) = 6 x^2 -x - 6 = 6 x^2 - x -12 = 0 (x-4)(x+3) = 0 x = 4 or -3